OpenStudy (anonymous):

Can you please help guide me through the last steps in this equation involving logarithms?

OpenStudy (anonymous):

\[1+\ln x = \ln4\]

OpenStudy (anonymous):

I am solving for x. Can you nudge me about the ln properties?

jimthompson5910 (jim_thompson5910):

1+ln(x) = ln(4) 1 = ln(4)-ln(x) 1 = ln(4/x) 4/x = e^1 4/x = e 4 = ex 4/e = x x = 4/e

OpenStudy (anonymous):

Thanks, I'll see if I can follow this on paper here... Just a sec...

jimthompson5910 (jim_thompson5910):

The properties I used were ln(x) - ln(y) = ln(x/y) and ln(x) = y converts to e^y = x

OpenStudy (anonymous):

I see the division coming into play with the subtraction.

OpenStudy (anonymous):

The last one is the key I was missing.

OpenStudy (anonymous):

you can also try \[\ln(x)=\ln(4)-1\] \[x=e^{\ln(4)-1}=e^{\ln(4)}\times e^{-1}=\frac{4}{e}\]

OpenStudy (anonymous):

ln x = e to what power gives us x Logarithms confuse me to no end, despite how "simple" they are supposed to be.

OpenStudy (anonymous):

I have duplicated the algebra, and it makes sense. Are you aware of a really good ln tutorial with exercises?

jimthompson5910 (jim_thompson5910):

try this page: http://www.purplemath.com/modules/solvelog.htm

OpenStudy (anonymous):

Thank you. If you answer that tutorial question in the new Q I posted, I'd be pleased to pin a medal upon you.